OpenStudy (anonymous):

What is the slope of the line between (3, –4) and (–2, 1)?

OpenStudy (zzr0ck3r):

(1-(-4))/(-2-3) = ?

OpenStudy (anonymous):

\[slope = \frac{ y2 - y1 }{x2 - x1 }\]

OpenStudy (anonymous):

I know the formula I just dont get what to do with the ordered pair at the end.

OpenStudy (zzr0ck3r):

it wont be an ordered pair

OpenStudy (anonymous):

There isn't a ordered pair at the end, you made a mistake somewhere.

OpenStudy (zzr0ck3r):

what do you get when you compute what I put(which I got from the fomula)?

OpenStudy (anonymous):

Try this remember two - equal a +\[\frac{ 1 +4 }{ -2 - 3 }\]

OpenStudy (anonymous):

The 1 -(-4) = 1 + 4

OpenStudy (anonymous):

you have to subtract on to[p and bottom...

OpenStudy (anonymous):

No solve the top and bottom seperate and keep it as a fraction: what is 1 + 4 over what is -2 - 3

OpenStudy (zzr0ck3r):

(1-(-4))/(-2-3) = (1+4)/-5 = 5/-5 = ?

OpenStudy (anonymous):

@barthelme1 Did you figure it out?

OpenStudy (anonymous):

1?

OpenStudy (zzr0ck3r):

thats 5/5 we did 5/-5

OpenStudy (zzr0ck3r):

note 5/-5 = -5/5

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

d(^_^)b

OpenStudy (anonymous):

Sweet